OpenStudy (anonymous):

Can someone please explain to me very elaborately ,in subtracting mixed fractions, how to solve the problem if the the first numerator is less than the next. For instance, numerators like 5 and 7 that can't be subtracted from each other. I'm having trouble understanding the borrowing process.

OpenStudy (amistre64):

realize that a number like:\[2\frac34=2+\frac34\]

OpenStudy (amistre64):

if its borrwing you need, borrow 1 do you ahve an example we can work?

OpenStudy (amistre64):

borrowing is just a way of rewriting a number into something equal but possibly simpler 523 = 500 + 20 + 3 -416 = -400 -10 - 6 523 = 500 + 20-10+10 + 3 -416 = -400 -10 - 6 523 = 500 + 10 +13 -416 = -400 -10 - 6 ------------------- 107 = 100 +0 +7

OpenStudy (amistre64):

we borrow a 'unit' measure to help out if need be. since a fraction is less than 1, the unit value to borrow is 1 \[~~~~2+\frac58\\-1-\frac78\] \[~~~~2-1+\frac88+\frac58\\-1-\frac78\] \[~~~~1+\frac{13}8\\-1-\frac78\\-------\\ ~~~~~0+\frac68 \]

OpenStudy (anonymous):

How about a question like...

OpenStudy (anonymous):

Understand that 9 is the same thing as 27/3, and that 7 is the same thing as 21/3

OpenStudy (anonymous):

Okay thanks! I got it.

OpenStudy (amistre64):

if you need to borrow, borrow 1 from the 9 and add it into the fraction (9-1) + (3/3 + 1/3) 8 + 4/3 -7 -2/3 --------- 1 +2/3